# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ IV-B4 Beyond-100 GHz: sub-THz and THz
Despite this potential, beyond-100 GHz faces severe constraints for NTNs: high free-space path loss, strong water-vapor absorption, weak diffraction, and hardware limits (low output power, thermal stability) restrict range and robustness [102, 103]. Consequently, sub-THz/THz should be viewed as complementary, beyond-5G options suited mainly to short-range, LoS inter-platform or backhaul links under controlled conditions—not for handheld service or long-range feeder links. Practical deployment will require multi-link coordination (e.g., fallbacks to FR2/FR3 or FSO), highly directive arrays with null steering, site/spatial diversity, and aggressive modulation/coding adaptation [95, 96].
IV-B5 FSO
@@
FSO technology is a valuable approach for future wireless networks, offering cost-effectiveness, easy deployment, high bandwidth, and strong security. The study in [104] provides an in-depth analysis of FSO systems, highlighting their role in next-generation optical communication. However, FSO propagation is challenged by atmospheric turbulence and alignment issues, necessitating mitigation strategies to maintain link stability. The study in [105] evaluates a ground-to-HAPS-to-RIS-to-UE communication system, where FSO is used for the ground-to-HAPS link, while RF communication is established for HAPS-to-RIS-to-UE transmissions. Key performance metrics, such as ergodic capacity, average bit error rate (BER), and outage probability, are analyzed using the bivariate Fox-H function. Results indicate that pointing errors and shadowing significantly affect system performance, with larger zenith angles and severe shadowing leading to degradation. Heterodyne detection improves efficiency by enhancing capacity and reducing BER and outage probability.
@@
The design of efficient HAPS-based NTNs requires balancing trade-offs between coverage, capacity, and resilience. A hybrid multi-band approach can enhance adaptability by leveraging different frequency bands for specific link requirements. For example, they suggest the following configurations:
  • Service links may operate at lower frequencies (sub-7 GHz, cmWave), to provide broad coverage and reliable signal propagation [75, 74, 82].
  • Feeder links may operate at cmWave and mmWave bands, to enable high-capacity, long-range transmission [106].
  • Inter-platform and backhaul links require ultra-high-speed data transfer, and may operate at sub-THz, THz, and FSO frequencies for seamless network integration.
This multi-band, hybrid spectrum strategy ensures a balance between coverage, capacity, and system resilience, supporting the development of robust HAPS-assisted NTN architectures.
IV-C 3GPP Specified Architectures
@@
As introduced earlier, in the 3GPP NTN framework [74], two primary platform payload configurations are defined: transparent (or bent-pipe) and regenerative. The choice of payload configuration determines the NTN architecture at a fundamental level, shaping its operational behavior. In the transparent architecture the platform acts as a transparent relay, forwarding signals directly without significant processing. A bent-pipe payload typically implements frequency conversion, analog filtering, and signal amplification. This straightforward design results in reduced payload complexity, lower power consumption on the platform, and generally lower development and deployment costs. However, the absence of onboard signal processing limits the system’s ability to handle dynamic traffic or execute advanced processing tasks. These functions are offloaded to the ground station, leading to a round-trip signal path (from the UE to the NTN platform and then to the ground station) which inherently increases latency.
@@
The regenerative architecture involves onboard signal processing, including tasks such as demodulation, decoding, and routing, in addition to the functions performed in the bent-pipe configuration. This onboard processing capability enables dynamic traffic management, such as load balancing based on QoS requirements. Platforms with sufficient computational resources can execute some UE tasks independently of the ground station, significantly reducing latency for time-sensitive applications. However, this capability comes at the cost of higher payload complexity and power consumption. The advanced technology required for regenerative systems also increases development and operational expenses.
@@
Both transparent and regenerative NTN configurations can be implemented in multiple ways, depending on the functionalities assigned to the platform. In a transparent (bent-pipe) architecture—used in 3GPP Rel-17 and Rel-18—the NTN platform acts as a radio repeater (with or without beam control) while the gNB remains on the ground; this limits payload complexity and enabled early deployment. Notably, the Rel-17 “Satellite Access Node” definition in 3GPP RAN4 still allowed some distribution of gNB functions between Earth and space while keeping RAN4 RF/performance specifications applicable. In contrast, Rel-19 introduces a regenerative (packet-processing) architecture in which a complete gNB is placed on the platform [107]. This enables native use of the $X_{n}$ interface over inter-satellite links (inter-gNB mobility, resource coordination, energy saving), on-payload packet switching and routing/handovers directly between satellites with minimal ground involvement, and reduced $U_{u}$ round-trip time (e.g., for random access and hybrid automatic repeat request (HARQ)). Regenerative payloads also provide the building block for Rel-19 “store-and-forward” (S&F) for IoT UEs, allowing continued operation if the feeder link is severed and even supporting UE-to-UE termination in space. These gains come at the expense of higher payload power and complexity. Earlier alternatives with only a gNB-distributed unit onboard (gNB-central unit on the ground) were found suboptimal due to heavy F1 signaling over the feeder link, lack of a standardized direct inter-satellite interface, and limited support for feeder-link switching and S&F. Consensus therefore converged on the full gNB-on-board approach.
@@
The role of MC in integrated TN-NTN systems depends on the operational objective. Beyond conventional dual- or multi-node connectivity, secondary MC nodes may be activated on demand to boost throughput. For example, [108] proposed a secondary-node activation algorithm, demonstrating a 9.1% increase in maximum average per-UE throughput compared to systems without MC, and a 5.3% improvement over baseline algorithms.
@@
In addition, MC enables dynamic traffic splitting between terrestrial and non-terrestrial nodes, allowing the network to balance load across layers and offload traffic from congested terrestrial nodes to NTN links. In [109], the authors introduced a framework for UE association, BS transmit power and activation, and bandwidth allocation between terrestrial and non-terrestrial tiers. Simulation results showed a 45% reduction in total power consumption and an approximately 250% increase in average throughput compared to standard 3GPP-based network models.
@@
Within integrated NTN systems, HAPS can assume diverse roles: from serving as terrestrial network extensions or satellite signal repeaters to operating as full-scale network centers equipped with baseband units and edge data processing capabilities. These roles directly impact latency, energy efficiency, and coverage performance. For instance, [110] proposed a quasi-stationary HAPS to enhance inter-satellite connectivity via FSO links with LEO satellites. Simulations under extreme conditions, such as volcanic activity, demonstrated the feasibility of supporting optical inter-satellite communication.
Experimental evaluations further confirm the versatility of HAPS-based systems. In [111], an automatic gain control system enabled a HAPS wireless repeater compatible with 5G NR to maintain stable transmission at maximum power, ensuring wide-area coverage. Similarly, [106] analyzed different HAPS-based link architectures and showed that HAPS can reduce latency in UAV ad-hoc networks, enhance energy efficiency in UAV cell-free architectures, improve coverage probability in satellite cell-free networks, and expand the service area of macro base stations through integrated access and backhaul designs.
@@
Nevertheless, real-world data highlight operational limitations. In [112], empirical analysis of Loon project flight tracking data revealed that while stratospheric balloons offer a cost-effective, delay-tolerant connectivity service, outages can last several hours, limiting the achievable service stability compared to planned terrestrial networks. Such constraints may explain the current trend toward fixed-wing HAPS designs in modern projects, which promise greater reliability and sustained performance.
IV-D HAPS-Based Network Trials and Deployments
@@
Over the past decades, numerous HAPS projects have explored applications in communication, as well as surveillance, disaster management, and environmental monitoring. While some achieved promising results, many faced technological, financial, or regulatory challenges that led to their termination even at advanced stages.
A prominent example is the Loon project [76], initiated by X Development LLC (formerly part of Google), which aimed to extend Internet connectivity to underserved regions using stratospheric balloons. Loon demonstrated HAPS potential by providing emergency connectivity in Peru (2017) during floods and in Puerto Rico after a hurricane, and by launching a commercial Internet service with Telkom Kenya in 2020 [112]. The balloons, following a regenerative architecture, carried an e-NodeB and antennas to serve ground UEs and communicate with gateways. Despite these achievements, concerns over commercial viability led to its discontinuation in 2021 [76].
@@
Zephyr, a fixed-wing solar-powered HAPS by Airbus [113], is a payload-agnostic platform providing persistent surveillance and communication. It supports both transparent and regenerative architectures, using either its in-house connectivity payload to extend 4G/5G networks or other compatible technologies, enabling platform-as-a-service. By combining communication, monitoring, and sensing functions, Zephyr offers opportunities for intelligent network optimization and QoS improvement.
Similarly, HAPSMobile, a collaboration between AeroVironment and SoftBank [77, 114], plans a network of Sungliders to provide connectivity to billions lacking wireless access and to support emerging 5G and IoT networks. SoftBank positions HAPS as “base stations in the sky” for wide-area coverage, recovery during large-scale disasters, and 3D connectivity for drones/UAVs in the 6G era, complementing its ongoing heavier-than-air development with lighter-than-air vehicles. No details found on the architecture. Pre-commercial stratospheric services are expected in Japan by 2026 [115].
V HAPS Technologies
@@
Technology achievements enable HAPS to enhance connectivity, optimize performance, and ensure sustainable operations. As an example, the authors in [2] explored the application of advanced techniques such as massive MIMO and mmWave communication into HAPS systems, identifying these as promising directions for future research. Similarly, the authors in [116] discussed key technological developments, trends, and challenges associated with HAPS. Nonetheless, this work primarily adopts an aeronautical engineering perspective, and lacks a thorough analysis from a telecommunications point of view. In addition, the study was published in April 2016, and does not reflect the most recent developments in the field. In this section, we address these limitations, and present an updated and comprehensive overview of recent technological trends that support the large-scale adoption of HAPS in the context of 6G networks. Specifically, we focus on the design of channel modeling (Sec. V-A), resource management (Sec. V-B), user assignment and cell association (Sec. V-C), multi-link connectivity (Sec. V-D), mobility management and handover (Sec. V-E), antenna (Sec. V-F), channel estimation (Sec. V-G), interference management and beamforming (Sec. V-H), energy efficiency (Sec. V-I), and ML (Sec. V-J) techniques for HAPS.
V-A Channel Modeling
Channel modeling for HAPS is evolving with new technologies, improving accuracy in predicting signal propagation, especially based on advanced simulation techniques and real-time data. Network simulators often lack realistic path loss and fading modules. To address this, models for rain, cloud, gaseous absorption, and Ricean fading have been developed [117]. Using QualNet, the authors showed how different attenuation types affect the packet delivery rate of HAPSground links, improving simulation realism and accuracy. The researchers in [118] used ray tracing simulations in suburban and urban environments, including downtown Chicago, to study HAPS-to-ground signal propagation. They developed a basic route loss model that estimates total propagation loss by considering terrestrial clutter and free-space propagation. The model adapts to different aerial vehicle heights and elevation angles using two environment-specific factors, and the study also analyzes multipath characteristics to support system design and fast fading modeling.
The authors of [119] analyzed a channel model to determine the capacity of HAPS-to- HAPS links, considering polarization mismatches, atmospheric gases, rain, and clouds/fog. Their simulations revealed the path gain characteristics and channel capacity across different antenna types, platform heights, and carrier frequency scenarios.
The study in [120] modeled HAPS-to-satellite communication channels with a focus on ionospheric effects. It showed that plasma density variations influence the channel while magnetic field effects on the refractive index can be neglected. Using the Appleton-Hartree model and an adapted Chapman model, the authors analyzed how Total Electron Content impacts path loss, supported by extensive numerical results.
A wideband channel measurement campaign at Ku- and X-bands investigated propagation effects in satellite-to-mobile radio links in four settings: urban, suburban, rural, and woodland. Results showed low coherence bandwidth in suburban and urban settings, suggesting channel equalization for high-data-rate services [121].

### v2

@@ V-C4 Beyond-100 GHz: sub-THz and THz
The beyond-100 GHz region is covering sub-THz typically defined between 100 and 300 GHz [31, 48] and even wider THz range (0.1–10 THz).It offers extreme bandwidth for immersive communications and sensing [184, 110]. Measurements at 140 GHz report 2.3 Gbps over 120 m, highlighting the rate–range trade-off and equalization needs [159]. Simulation studies also show that integrated sensing and communication (ISAC) in these bands can outperform 5G-spectrum baselines [119]. For NTN topologies, sub-THz has been proposed for backhaul between UAV-BSs and CPUs, with HAPS acting as the CPU to preserve LoS [14]. Feasibility analyses indicate that THz feeder/backhaul links around 10 Gbps are possible in light rainfall given careful link budgets and signal processing [206] Reconfigurable intelligent surfaces (RISs) can further alleviate blockage [26]. Despite this potential, beyond-100 GHz faces severe constraints for NTNs: high free-space path loss, strong water-vapor absorption, weak diffraction, and hardware limits (low output power, thermal stability) restrict range and robustness [33, 200]. Consequently, sub-THz/THz should be viewed as complementary, beyond-5G options suited mainly to short-range, LoS inter-platform or backhaul links under controlled conditions—not for handheld service or long-range feeder links. Practical deployment will require multi-link coordination (e.g., fallbacks to FR2/FR3 or FSO), highly directive arrays with null steering, site/spatial diversity, and aggressive modulation/coding adaptation [184, 110].
V-C5 FSO
@@
FSO frequency region covering an ultra-wide range of spectrum, roughly, between 3 THz and 30 PHz, which include infrared, visible light, and ultraviolet categories and basically overlap with THz region. It is worth noting, however, that unlike RF bands, no single bounded range is allocated for FSO by regulations, and discrepancies in the reported values may therefore occur. FSO systems offers a broad range of indoor and outdoor applications including cyber security, surveillance and monitoring, inter-building communication, high-capacity backhauling, etc. [106].FSO technology is a valuable approach for future wireless networks, offering cost-effectiveness, easy deployment, high bandwidth, and strong security. Samir et al. [17] provided an in-depth analysis of FSO systems, highlighting their role in next-generation optical communication. However, FSO propagation is challenged by atmospheric turbulence and alignment issues, necessitating mitigation strategies to maintain link stability. Shang et al. [194] evaluated a ground-to-HAPS-to-RIS-to-UE communication system, where FSO is used for the ground-to-HAPS link, while RF communication is established for HAPS-to-RIS-to-UE transmissions. Key performance metrics, such as ergodic capacity, average bit error rate (BER), and outage probability, are analyzed using the bivariate Fox-H function. Results indicate that pointing errors and shadowing significantly affect system performance, with larger zenith angles and severe shadowing leading to degradation. Heterodyne detection improves efficiency by enhancing capacity and reducing BER and outage probability.
@@
The design of efficient HAPS-based NTNs requires balancing trade-offs between coverage, capacity, and resilience. A hybrid multi-band approach can enhance adaptability by leveraging different frequency bands for specific link requirements. For example, (i) service links may operate at lower frequencies (sub-7 GHz, cmWave), to provide broad coverage and reliable signal propagation [4, 2, 1]; (ii) feeder links may operate at cmWave and mmWave bands, to enable high-capacity, long-range transmission [142]; (iii) Inter-platform and backhaul links require ultra-high-speed data transfer, and may operate at sub-THz, THz, and FSO frequencies for seamless network integration. This multi-band, hybrid spectrum strategy ensures a balance between coverage, capacity, and system resilience, supporting the development of robust HAPS-assisted NTN architectures.
Figure 3: Transparent NTN architecture.
V-D 3GPP Specified Architectures
@@
As introduced earlier, in the 3GPP NTN framework [2], two primary platform payload configurations are defined: transparent (or bent-pipe) and regenerative. The choice of payload configuration determines the NTN architecture at a fundamental level, shaping its operational behavior. In the transparent architecture (see Fig. 3), the platform acts as a transparent relay, forwarding signals directly without significant processing. A bent-pipe payload typically implements frequency conversion, analog filtering, and signal amplification. This straightforward design results in reduced payload complexity, lower power consumption on the platform, and generally lower development and deployment costs. However, the absence of onboard signal processing limits the system’s ability to handle dynamic traffic or execute advanced processing tasks. These functions are offloaded to the ground station, leading to a round-trip signal path (from the UE to the NTN platform and then to the ground station) which inherently increases latency.
@@
The regenerative architecture (see Fig. 4) involves onboard signal processing, including tasks such as demodulation, decoding, and routing, in addition to the functions performed in the bent-pipe configuration. This onboard processing capability enables dynamic traffic management, such as load balancing based on QoS requirements. Platforms with sufficient computational resources can execute some UE tasks independently of the ground station, significantly reducing latency for time-sensitive applications. However, this capability comes at the cost of higher payload complexity and power consumption. The advanced technology required for regenerative systems also increases development and operational expenses.
@@
Both transparent and regenerative NTN configurations can be implemented in multiple ways, depending on the functionalities assigned to the platform. In a transparent (bent-pipe) architecture—used in 3GPP Rel-17 and Rel-18—the NTN platform acts as a radio repeater (with or without beam control) while the gNB remains on the ground; this limits payload complexity and enabled early deployment. Notably, the Rel-17 “Satellite Access Node” definition in 3GPP RAN4 still allowed some distribution of gNB functions between Earth and space while keeping RAN4 RF/performance specifications applicable. In contrast, Rel-19 introduces a regenerative (packet-processing) architecture in which a complete gNB is placed on the platform [232]. This enables native use of the $X_{n}$ interface over inter-satellite links (inter-gNB mobility, resource coordination, energy saving), on-payload packet switching and routing/handovers directly between satellites with minimal ground involvement, and reduced $U_{u}$ round-trip time (e.g., for random access and hybrid automatic repeat request). Regenerative payloads also provide the building block for Rel-19 “store-and-forward” (S&F) for IoT UEs, allowing continued operation if the feeder link is severed and even supporting UE-to-UE termination in space. These gains come at the expense of higher payload power and complexity. Earlier alternatives with only a gNB-distributed unit onboard (gNB-central unit on the ground) were found suboptimal due to heavy F1 signaling over the feeder link, lack of a standardized direct inter-satellite interface, and limited support for feeder-link switching and S&F. Consensus therefore converged on the full gNB-on-board approach.
Figure 4: Regenerative NTN architecture.
@@
The role of MC in integrated TN-NTN systems depends on the operational objective. Beyond conventional dual- or multi-node connectivity, secondary MC nodes may be activated on demand to boost throughput. For example, Majamaa, Martikainen, Sormunen, and Puttonen [149] proposed a secondary-node activation algorithm, demonstrating a 9.1% increase in maximum average per-UE throughput compared to systems without MC, and a 5.3% improvement over baseline algorithms.
@@
In addition, MC enables dynamic traffic splitting between terrestrial and non-terrestrial nodes, allowing the network to balance load across layers and offload traffic from congested terrestrial nodes to NTN links. Alam et al. [18] introduced a framework for UE association, BS transmit power and activation, and bandwidth allocation between terrestrial and non-terrestrial tiers. Simulation results showed a 45% reduction in total power consumption and an approximately 250% increase in average throughput compared to standard 3GPP-based network models.
@@
Within integrated NTN systems, HAPS can assume diverse roles: from serving as terrestrial network extensions or satellite signal repeaters to operating as full-scale network centers equipped with baseband units and edge data processing capabilities. These roles directly impact latency, energy efficiency, and coverage performance. For instance, Erdogan et al. [65] proposed a quasi-stationary HAPS to enhance inter-satellite connectivity via FSO links with LEO satellites. Simulations under extreme conditions, such as volcanic activity, demonstrated the feasibility of supporting optical inter-satellite communication. Experimental evaluations further confirm the versatility of HAPS-based systems. Hasegawa, Konishi, Ohta, and Nagate [89] used an automatic gain control system, which enabled a HAPS wireless repeater compatible with 5G NR to maintain stable transmission at maximum power, ensuring wide-area coverage. Similarly, Lou et al. [142] analyzed different HAPS-based link architectures and showed that HAPS can reduce latency in UAV ad-hoc networks, enhance energy efficiency in UAV cell-free architectures, improve coverage probability in satellite cell-free networks, and expand the service area of macro base stations through integrated access and backhaul designs.
@@
Nevertheless, real-world data highlight operational limitations. Empirical analysis of Loon project flight tracking data revealed that while stratospheric balloons offer a cost-effective, delay-tolerant connectivity service, outages can last several hours, limiting the achievable service stability compared to planned terrestrial networks. Such constraints may explain the current trend toward fixed-wing HAPS designs in modern projects, which promise greater reliability and sustained performance [189].
V-E HAPS-Based Projects
@@
Over the past decades, numerous projects have explored various applications of HAPS, particularly in communication, but also in fields like surveillance, disaster management, and environmental monitoring. While some of these projects achieved promising results, many encountered significant challenges that led to their closure, even at advanced stages of development. These challenges included technological difficulties, unsustainable financial models, or regulatory hurdles.
One notable example is the Loon project [229], initiated by X Development LLC, formerly part of Google’s innovation-focused division. The project aimed at extending Internet connectivity to underserved and remote areas using stratospheric balloons. Loon achieved several milestones, showcasing the potential of HAPS-based solutions. For instance, in 2017, the team successfully navigated balloons to disaster-stricken regions, providing emergency connectivity in Peru during severe flooding and in Puerto Rico after a hurricane destroyed critical ground infrastructure. In 2020, Loon partnered with Telkom Kenya to launch a commercial Internet-via-balloon service for Telkom Kenya subscribers, marking a significant step toward scalable deployment. The balloons, following a regenerative architecture, were equipped with an e-NodeB and various antennas to transmit LTE signals to ground UEs and communicate with each other and with gateway stations [189].
However, despite an important role and a number of delivered breakthroughs in both academia and industry, the Loon project was ended in 2021. Its shutdown stemmed from many reasons including economical (high capital expenditure and operating costs), regulatory (spectrum allocation restrictions, international airspace coordination and data sovereignty issues), technological (channel modeling and network planning challenges, complicated station-keeping) and environmental/resource (helium scarcity, power limitations from photo-voltaic panels, occasional environmental pollution caused by polyethylene balloon wreckage) [244].
Another ambitious project, Aquila, initiated by Facebook in 2014, aimed to develop a HAPS system based on solar-powered unmanned aircraft to deliver broadband connectivity to underserved and remote regions. Operating in the stratosphere, Aquila was designed as a quasi-stationary relay providing line-of-sight coverage over areas of several thousand square kilometers, complementing terrestrial networks where infrastructure deployment is economically unviable. The project emphasized long-endurance flights (target: 60–90 days), solar energy harvesting, and mmWave communication payloads for high-capacity air-to-ground links. Despite promising early demonstrations and due to difficulties with in-house aerospace expertise led to insufficient solar energy harvesting and battery efficiency, inadequate flight endurance and, above all, structural integrity issues, their own aircraft development was ended in 2018 in favor of ecosystem partnerships with aerospace firms such as Airbus [87].
@@
Zephyr, a series of HAPS aircraft initially designed by Qinetiq and produced by Airbus, effectively became a successor of Aquila [171]. Now it is an ongoing fixed-wing solar-powered HAPS project by Airbus, is designed as a payload-agnostic platform capable of extended operation in the stratosphere [10]. It provides persistent surveillance and communication services with various types of payloads. Zephyr supports both transparent and regenerative network architectures, utilizing either its default in-house connectivity payload designed to extend 4G and 5G networks, or by integrating other compatible technologies, thereby functioning as platform-as-a-service. Combining a wide range of monitoring, tracking, sensing and detection applications with a communication payload, this platform can offer great opportunities for intelligent network optimization and QoS improvement.
Similarly, HAPSMobile, a collaborative project between AeroVironment and SoftBank, aims at establishing a network of Sungliders [198] to provide connectivity to billions of people lacking basic wireless communication, and to supply additional bandwidth for emerging 5G and IoT networks [16]. SoftBank also announced that, in partnership with the U.S.-based company Sceye Inc., it plans to launch pre-commercial HAPS stratospheric telecommunications services in Japan in 2026. [199]. In these systems, where the HAPS acts as a flying BS, a regenerative payload architecture is the likely choice, although details are hard to find at this stage.
Failures of the Loon and Aquila projects highlighted several major lessons that must be learned to enable more successful future development of HAPS-based solutions. From a technological perspective, it is crucial to ensure platform maturity through iterative prototyping and proper alignment of expertise. Aquila’s structural integrity issues and limited endurance stemmed from insufficient aerospace-domain knowledge. It is also important to consider economic sustainability in addition to technical feasibility. Technological proof-of-concepts (e.g., Loon’s over 300-day flights or Aquila’s mmWave links) are insufficient without a viable business model. Early-stage techno-economic analyses, including sensitivity to subsidies or public–private partnerships, are required to avoid over-optimistic assumptions. Similarly, environmental and resource constraints must be accounted for in the design phase. Helium scarcity limited Loon’s scalability, while Aquila’s solar–battery efficiency proved inadequate under variable conditions. Consequently, it is essential to adopt sustainable alternatives aligned with 6G sustainability goals, emphasizing energy-efficient NTN solutions to reduce ecological footprints. Regulatory foresight and spectrum allocation strategies should also be embedded into project roadmaps and aligned with International Telecommunication Union (ITU) and 3GPP frameworks in order to avoid conflicts with local regulations and international coordination challenges. These issues may be mitigated through closer academia–industry collaboration and effective partnerships across multiple industrial sectors, including aerospace, communications, energy harvesting, and materials engineering. For this purpose, future HAPS research should prioritize system-level prototyping using open-source tools, demand-driven deployment strategies, and technological cooperation. As an example of such collaboration, the HAPS Alliance, founded in 2020, brings together globally recognized companies from telecommunications, technology, aviation, and aerospace, as well as academic research groups. Moreover, shifting from competition for market share among terrestrial providers, SatCom operators, and emerging HAPS-based projects toward mutually beneficial collaboration may represent a key step toward a fully integrated TN-NTN architecture incorporating multiple layers to achieve truly ubiquitous and seamless connectivity.
VI HAPS Technologies
@@
Technology achievements enable HAPS to enhance connectivity, optimize performance, and ensure sustainable operations. As an example, Karabulut Kurt et al. [116] explored the application of advanced techniques such as massive MIMO and mmWave communication into HAPS systems, identifying these as promising directions for future research. Similarly, d’Oliveira, Melo, and Devezas [55] discussed key technological developments, trends, and challenges associated with HAPS. Nonetheless, this work primarily adopts an aeronautical engineering perspective, and lacks a thorough analysis from a telecommunications point of view. In addition, the study was published in April 2016, and does not reflect the most recent developments in the field.
In this section, we address these limitations, and present an updated and comprehensive overview of recent technological trends that support the large-scale adoption of HAPS in the context of 6G networks. We focus on key enablers such as channel modeling (Sec. VI-A), resource management (Sec. VI-B), user assignment and cell association (Sec. VI-C), multi-link connectivity (Sec. VI-D), mobility management and handover (Sec. VI-E), antenna (Sec. VI-F), channel estimation (Sec. VI-G), interference management and beamforming (Sec. VI-H), energy efficiency (Sec. VI-I), ML (Sec. VI-J) and Cybersecurity (Sec. VI-K) techniques for HAPS (See Table LABEL:tab:haps_technologies_summary in Appendix). In order to review technology achievements systematically we classify studies using a multi-dimensional taxonomy based on following dimensions: research methodology (analytical, simulation-based, experimental); network integration level (standalone HAPS, HAPS-TN, HAPS-NTN, HAPS-TN-NTN, no HAPS); and related frequency bands (Sub-7 GHz, mmWave/FR2, Sub-THz/THz, and FSO). From Table LABEL:tab:haps_technologies_summary, it can be observed that analytical research methodology dominates the body of HAPS-related studies. This prevalence reflects the current maturity level of HAPS technology, where theoretical frameworks enable rapid feasibility assessments, while simulation-based investigations are limited by the availability of validated channel and system models. Experimental studies remain scarce due to regulatory, financial, and technical barriers. Consequently, researchers often rely on self-developed simulation models with limited capabilities in order to validate their analytical findings. Nevertheless, collaborative initiatives—such as those led by the HAPS Alliance—may facilitate the collection of empirical data and contribute to the standardization of datasets for more reliable and reproducible simulation studies. It is also worth noting that the majority of existing studies focus on the sub-7 GHz and mmWave/FR2 frequency ranges, while the emerging cmWave and THz bands remain comparatively underexplored.
VI-A Channel Modeling
Channel modeling for HAPS is evolving with new technologies, improving accuracy in predicting signal propagation, especially based on advanced simulation techniques and real-time data. However, existing HAPS-related channel modeling studies predominantly focus on HAPS-to-ground links, whereas inter-platform channel characteristics remain comparatively underrepresented.
Network simulators often lack realistic path loss and fading modules. To address this, models for rain, cloud, gaseous absorption, and Ricean fading have been developed. Using QualNet, Zheng, Wang, and Meng [251] showed how different attenuation types affect the packet delivery rate of HAPS-to-ground links on 10, 15, 20 and 25 GHz communication frequency, improving simulation realism and accuracy. Also, Zain et al. [241] showed Ka-band rain attenuation is about 10× higher than in C-band and strongly location-dependent, and presented an analytical model based on the ITU-R empirical model to predict Earth–space attenuation for fully and partially rain-affected paths. For a 1-km cell, attenuation reaches 43 dB at 0.01% availability and roughly doubles as cell size increases.
To address the complexity of propagation models in suburban and urban environments, Hsieh et al. [97] applied ray tracing simulations to study HAPS-to-ground signal propagation in Chicago downtown. They developed a path loss model in order to estimate total propagation loss by considering terrestrial clutter and free-space propagation. The model adapts to different aerial vehicle heights and elevation angles using two environment-specific factors, and the study also analyzes multipath characteristics to support system design and fast fading modeling. To examine the inter-platform propagation, Yılmaz et al. [237] analyzed a channel model of HAPS-to-HAPS links in order to determine its capacity while considering polarization mismatches, atmospheric gases, rain, and clouds/fog. Their simulations revealed the path gain characteristics and channel capacity across different antenna types, platform heights, and carrier frequency scenarios. Meanwhile Palanci et al. [170] modeled HAPS-to-satellite communication channels with a focus on ionospheric effects. It showed that plasma density variations influence the channel while magnetic field effects on the refractive index can be neglected. Using the Appleton-Hartree model and an adapted Chapman model, the authors analyzed how Total Electron Content impacts path loss, supported by extensive numerical results.